OpenStudy (anonymous):

Can someone please help me I'm lost Solve for x: −5|x + 1| = 10 x = 0 x = −3 and x = 1 x = −1 and x = 3 No solution

OpenStudy (solomonzelman):

Divide both sides by `-5` and you will see that there is no solution to the equation, because you are going to get the absolute value of something equal to a negative, which is not possible.

OpenStudy (anonymous):

So it would be no solution

OpenStudy (solomonzelman):

Yup

OpenStudy (anonymous):

Thankyou, is it because it was a negative and a positive?

OpenStudy (solomonzelman):

I don't really get what you mean by `it was a negative and a positive`

OpenStudy (anonymous):

Um because when I was checking my solution I got, -10=10 and 10=10

OpenStudy (solomonzelman):

Well... \(\normalsize\color{ purple}{ −5|x + 1| = 10}\) divide both sides by `-5` and you get, \(\normalsize\color{ purple}{ |x + 1| = -2}\)

OpenStudy (solomonzelman):

And this is where there is no solution, because \(\normalsize\color{ purple}{ |\rm{~anything~}| ≥ 0}\)

OpenStudy (solomonzelman):

Makes sense, right ?

OpenStudy (anonymous):

Ok I get it now, thank you :)

OpenStudy (solomonzelman):

Anytime :)

OpenStudy (anonymous):

Are you available for more questions?

OpenStudy (solomonzelman):

Yes, perhaps.

OpenStudy (anonymous):

Well standby, because I'm so confused on this absolute value

OpenStudy (anonymous):

The minimum and maximum temperature on a cold day in Lollypop Town can be modeled by the equation below: 2|x − 6| + 14 = 38 What are the minimum and maximum temperatures for this day? x = −9, x = 21 x = −6, x = 18 x = 6, x = −18 No solution

OpenStudy (anonymous):

Im lost to

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 2|x − 6| + 14 = 38 }\) \(\normalsize\color{blue}{ 2|x − 6| + 14\color{red}{ -14 } = 38\color{red}{ -14 } }\) \(\normalsize\color{blue}{ 2|x − 6| = 24 }\) \(\normalsize\color{blue}{ 2|x − 6|\color{darkgoldenrod}{ \div2 } = 24 \color{darkgoldenrod}{ \div2 } }\) \(\normalsize\color{blue}{ |x − 6| = 12 }\) \(\normalsize\color{blue}{ |x − 6| = 12 ~~~~\LARGE\color{black}{ ⇒ }~~~~}\)\(\normalsize\color{blue}{ x − 6 = 12 ~~~~~~~\rm{or}~~~~~~x-6=-12}\)

OpenStudy (solomonzelman):

So, what is the maximum, and what is the minimum ? Let me know if you need more help.

OpenStudy (xapproachesinfinity):

well first subtract 14 from both sides then divide by 2 both sides

OpenStudy (anonymous):

So No solution?

OpenStudy (solomonzelman):

According to my post there is a solution

OpenStudy (anonymous):

Wouldnt it be C?

OpenStudy (solomonzelman):

Close, but the 6 is not positive, and the 18 in not negative.

OpenStudy (anonymous):

Ok so B. Thank you so much

OpenStudy (solomonzelman):

Anytime

OpenStudy (anonymous):

Solve for x: |x − 2| + 10 = 12 x = 0 and x = 4 x = −4 and x = 0 x = −20 and x = 4 No solution

OpenStudy (anonymous):

I got close to the end, then got confused

OpenStudy (solomonzelman):

show me what you got, please

OpenStudy (anonymous):

I chose C. Because I substituted

OpenStudy (anonymous):

@solomonzelman

OpenStudy (solomonzelman):

I will show the steps, and you will do the work, okay ?

OpenStudy (anonymous):

Ok :)

OpenStudy (solomonzelman):

1) Subtract `10` from both sides 2) when you have \(\normalsize\color{blue}{ │\rm{something}│ =number }\) we get, \(\normalsize\color{blue}{ \rm{something} =number ~~~~~or~~~~\rm{something} =-~number }\) Appply this here.

OpenStudy (solomonzelman):

questions ?

OpenStudy (anonymous):

Ok, so it wouldn't be a solution, because they wouldnt equal